What is the present value of $90,000 due 7 years from now, discounted at 9%? (For calculation purposes, use 5 decimal places as displayed in the factor table provided. Round answer to 2 decimal places, e.g. 52.75.)

Present value $

What is the present value of $150,000 due 5 years from now, discounted at 12%? (For calculation purposes, use 5 decimal places as displayed in the factor table provided. Round answer to 2 decimal places, e.g. 52.75.)

Present value $
Answer #1

Present value=Cash flows*Present value of discounting factor(rate%,time period)

a.Present value=90,000/1.09^7

=90,000*0.54703

=$49232.7

b.Present value=150,000/1.12^5

=150,000*0.56743

=$85114.5